OpenStudy (anonymous):

Find the volume of a cone with a radius of 6 m and a height of 10 m.

OpenStudy (anonymous):

1/3pir^2h

OpenStudy (studygurl14):

\[V = \frac{ 1 }{ 3 }\pi r^{2}h\] where r = radius and h=height so.... \[V = \frac{ 1 }{ 3 }\pi (6)^{2}(10)\] can you do the rest?

OpenStudy (anonymous):

yes

OpenStudy (studygurl14):

Good luck!